OpenStudy (anonymous):

What is the slope of the line passing through the points (0, −5) and (4, 2)?

OpenStudy (anonymous):

@vzfreakz @just_one_last_goodbye @Catseyeglint911 @GracieBugg @GracieBugg @Owlcoffee @Babynini @KAKES1967 @Question_Asker @xisaaaiahx

OpenStudy (anonymous):

The slope is 7/4

OpenStudy (anonymous):

tysm

OpenStudy (anonymous):

What is the slope of the line whose equation is 2x – 4y = 10?

OpenStudy (anonymous):

can you help me w that one also

OpenStudy (anonymous):

@TheSavior

OpenStudy (lanhikari22):

Just change the form of the equation to look like y = mx+b. m is your slope. You have 2x - 4y = 10, solve for y, and see what your slope is.

OpenStudy (lanhikari22):

slope is m

OpenStudy (anonymous):

idk how can you tell me what it is @LanHikari22

OpenStudy (lanhikari22):

I can't. you apparently get banned when you just tell answers here

OpenStudy (lanhikari22):

also to solve for something you just take everything else to the other side and leave y alone in one side. Like, y - x = 0, you add x in both sides to get y = x. 5y + 2x = 0. you subtract 2x in both sides to get rid of 2x in the left side, and you get 5y = -2x. Now you divide both sides by 5 to get rid of the five. Just do it so that you can get rid of everything in the left side except for y.

OpenStudy (lanhikari22):

solve your problem and the number attached to your x is your slope. y = mx+b

OpenStudy (babynini):

Do you get it? :) there is a formula for this! do you know it?